Without graphing determine whether the following pairs of lines are (a) parallel (b) perpendicular, or (c) neither parallel nor perpendicular 5x-6y=19
6x+5y=-30

Answer by Alan3354(69443)

5x-6y=19 m1 = 5/6
6x+5y=-30 m2 = -6/5
m1 * m2 = -1 --> perpendicular